1. A random sample of n = 60 measurements is obtained from a population with μ= 192 and σ = 43. Describe the sampling distribution for the sample means by computing the μx and  σx.
1
Expert's answer
2021-03-30T07:32:50-0400

Solution:

Given, "n=60, \mu=192,\sigma=43"

Mean of 60 sample means, "\mu_{x}=\mu=192"

Standard deviation of 60 sample means, "\sigma_x=\dfrac{\sigma}{\sqrt n}=\dfrac{43}{\sqrt {60}}\approx5.551"

Thus, sampling distribution has "\mu_{x}=192,\sigma_x=5.551"